Calculate the standard molar entropies at 25 oC for the following reactions.

2CH_4(g) + O_2(g) right arrow 2 CH_3OH(I) 4 Al(s)
Answer #1

a)

Srxn = Sproduct - Sreactant

Srxn = 2*SCH3OH - (2*SCH4 + SO2)

Srxn = 2*(126.8 ) - (2*186.2 +205.0 ) = - 323.8 J/molK

b)

Srxn = Sproduct - Sreactant

Srxn = 2*SAl2O3 - (4SAl + 3¨SO2)

Srxn = 2*50.9 - (4*28.3 l + 3*205.0 ) = -626.4 J/molK